# Deep Reinforcement Learning for Dynamic Order Picking in Warehouse Operations

### v1

@@ 4 Methodology
In this section, we present the formulation of the real-time warehouse order picking problem. We model this problem using a finite Markov Decision Process (MDP) with discrete time steps to capture its dynamic nature. Specifically, each time step $t$ involves making a movement or a decision regarding the picker’s actions. At each step, the system state $S_{t}$ is observed, encapsulating information about both the orders, their locations in warehouse and the picker. Based on this state, an action $a_{t}$ is chosen, determining the picker’s movement direction, whether the picker should remain stationary, or if items should be released at the depot. Executing this action leads to a new system state $S_{t+1}$, from which the next action $a_{t+1}$ is determined for the subsequent time step $t+1$.
@@ 4.1 MDP Formulation
    State: The system state at time step $t$ is denoted as $\boldsymbol{S_{t}}=(\boldsymbol{S^{p}_{t}},\boldsymbol{S^{o}_{t}})$, where $\boldsymbol{S^{p}_{t}}$ represents the picker state and $\boldsymbol{S^{o}_{t}}$ represents the state of existing orders in the warehouse.
@@
    The picker state, $\boldsymbol{S^{p}_{t}}=(S^{H}_{t},S^{V_{1}}_{t},S^{V_{2}}_{t},S^{C}_{t})$, consists of four components: $S^{H}_{t}$, $S^{V_{1}}_{t}$, $S^{V_{2}}_{t}$, and $S^{C}_{t}$. Here, $S^{H}_{t}$ represents the horizontal position of the picker, which is the relative position of the picker with respect to cross-aisles, $S^{V_{1}}_{t}$ and $S^{V_{2}}_{t}$ represent the vertical position, which are picker’s positions with respect to pick aisles, and $S^{C}_{t}$ indicates the remaining capacity of the picker. If $S^{C}_{t}=0$, the picker is full and cannot pick up additional items; if $S^{C}_{t}=K$, the picker is empty. The horizontal state $S^{H}_{t}$ can take values of -1, 0, or 1: -1 indicates the picker is at the back cross-aisle, 1 indicates the front cross-aisle, and 0 indicates the picker is within one of the pick aisles. The exact location of the picker within a pick aisle when $S^{H}_{t}=0$ is implicitly encoded in the order state $\boldsymbol{S^{o}_{t}}$. The vertical position is determined by $S^{V_{1}}_{t}$ and $S^{V_{2}}_{t}$, which equal $12n-$ 1 and $2n$ when the picker is at aisle $n$. The rationale for representing each pick aisle by two consecutive indices, rather than a single index, lies in the formulation of $\boldsymbol{S^{o}_{t}}$ as a vector of size $2N$. Enhancing this consistency between the two components of the state helps in the neural network’s interpretation.
@@
    The state of orders, $\boldsymbol{S^{o}_{t}}=(S^{1}_{t},S^{2}_{t},\ldots,S^{2N-1}_{t},S^{2N}_{t})$, where $N$ is the number of pick aisles in the warehouse. Each aisle $n$ is represented by $S^{2n-1}_{t}$ and $S^{2n}_{t}$. These values indicate the state of orders in each pick aisle depending on the picker’s vertical movement direction. $S^{2n-1}_{t}$ represents the value of orders if the picker moves upward in its current aisle, while $S^{2n}_{t}$ represents the value if the picker moves downward. These values are computed by summing the value of each order in the aisle divided by its distance from the picker, based on the direction of movement. The distance between the picker and an order is the smallest number of movements required for picker to reach the order in terms of the number of location storage. These states are calculated as
    $S^{2n-1}_{t}=\displaystyle\sum_{i=1}^{L}\displaystyle\frac{n_{i}}{\rho^{1}_{i}}\mbox{ and }S^{2n}_{t}=\displaystyle\sum_{i=1}^{L}\displaystyle\frac{n_{i}}{\rho^{2}_{i}},$
@@
    This formulation implicitly captures the importance of the picker’s visiting direction for each aisle. Notably, when $S^{H}_{t}=-1$, then $S^{2n}_{t}=0$ for all aisles since the picker cannot move upward. Similarly, when $S^{H}_{t}=1$, then $S^{2n-1}_{t}=0$ for all aisles. The starting state is $\boldsymbol{S^{o}_{0}}=\boldsymbol{0}$ since there are no orders at time zero, and $\boldsymbol{S^{p}_{t}}=(1,2d-1,2d,K)$, where $d$ is the aisle where depot is located at, and $K$ is the picker capacity.
@@
    Action: Given the state $S_{t}$, the action $a_{t}$ determines the picker’s decision. The action $a_{t}$ can take on values from 0 to 4. Specifically, $a_{t}=0$ signifies that the picker drops off items if located at the depot, and otherwise results in no movement. Actions with values 1 and 2 correspond to moving the picker to the right and left, respectively. In this study, actions are constrained to feasible options only. For instance, when the picker is situated within a pick aisle (i.e., $S^{H}_{t}=0$), actions 1 and 2 are disallowed. Actions 3 and 4 correspond to moving the picker upward and downward, respectively. Additionally, when the picker is in the back cross-aisle (i.e., $S^{H}_{t}=-1$), action 3 is prohibited, and when the picker is in the front cross-aisle (i.e., $S^{H}_{t}=1$), action 4 is prohibited.
@@
    $r_{t}=\begin{cases}-1&\text{if }a_{t}=0\text{ and picker not at the depot}\\ R\cdot(K-S^{C}_{t})\cdot\alpha&\text{if }a_{t}=0\text{ and picker at the depot}\\ -n^{m}_{t}+R\cdot n^{p}_{t}&\text{otherwise}\end{cases}$
    Here, $R$ denotes the reward for picking up an item, $n^{m}_{t}$ represents the number of storage locations the picker moves during the time step, and $n^{p}_{t}$ indicates the number of items picked up by the picker. When the picker is at the depot location and $a_{t}=0$, this action implies unloading all items. The number of items in the picker is $K-S^{C}_{t}$, and $\alpha$ is a coefficient that modulates the reward for unloading items relative to picking them up. The value of $\alpha$ ranges between 0 and 1, where smaller values signify that picking up items is more valuable than unloading them at the depot, while larger values imply equal value for both actions. This parameter is designed to balance the trade-off between average time and average distance per order. It is anticipated that with smaller values of $\alpha$, the picker would attempt to maximize its capacity usage before heading towards the depot, thereby minimizing traveling distance. Conversely, with larger values of $\alpha$, the picker would prefer to unload items as soon as possible, minimizing both wait time and delivery time.
@@ 4.2 Proposed Approach
Our objective is to train a policy that aims to maximize the discounted cumulative reward $R_{t}=\sum_{k=0}^{\infty}\gamma^{k}\cdot r_{t+k}$, where $R_{t}$ is referred to as the return, and $10<\gamma<$ 1 is the discount factor that balances the importance between immediate and future rewards. Small values of $\gamma$ indicate that immediate actions are more significant than future actions, while larger values of $\gamma$ place greater emphasis on future rewards.
The fundamental concept behind Q-learning is that if we had access to a function $Q^{*}:\text{State}\times\text{Action}\rightarrow\mathbb{R}$, which could inform us of the expected return if we were to take a particular action in a given state, we could straightforwardly construct a policy that maximizes our rewards: $\pi^{*}=\arg\max_{a}Q^{*}(s,a)$, where $Q^{*}(s,a)=\max(E_{\pi^{*}}[R_{t}\mid s_{t}=s,a_{t}=a])$. However, due to the complexity of the environment, we do not have direct access to $Q^{*}$. Nevertheless, since neural networks are universal function approximators, we can construct a neural network and train it to approximate $Q^{*}$.
Figure 2: Proposed Deep Neural Network Architecture
@@ 4.2.1 Architecture of Deep Neural Network
The deep neural network architecture designed for this study is specifically tailored to handle the state representation and predict optimal actions for the picker in a dynamic warehouse environment. The network serves as a Q-network, approximating the action-value function $Q(s,a)$, and consists of several fully connected layers to process both the picker state and the order state separately before combining them for the final action prediction. Figure 2 provides a summary of the architecture.
The architecture is structured as follows:
  • Input Layers: The input to the neural network is divided into two parts: the picker state $S^{p}_{t}$ and the order state $S^{o}_{t}$. The picker state $S^{p}_{t}$ is a four-dimensional vector represented as $(S^{H}_{t},S^{V_{1}}_{t},S^{V_{2}}_{t},S^{C}_{t})$, while the order state $S^{o}_{t}$ is a vector of size $2N$, where $N$ is the number of aisles in the warehouse.
@@
This deep neural network architecture effectively captures the complex interactions between the picker and the orders in the warehouse, enabling it to predict optimal actions that maximize the cumulative reward over time.
@@ 4.2.2 Training of Deep Neural Network
The training procedure for each mini-batch is a multi-step process. First, the states, next states, actions, and rewards within the batch are stacked for parallelization. Then, the policy network is set to training mode, while the target network enters evaluation mode. This setup allows for the computation of state-action values using the policy network, and expected state-action values using the target network. The discrepancy between these values is then used to calculate the loss, and we utilize the Huber loss function for robustness. The calculated loss is then backpropagated through the policy network, facilitating its optimization. To ensure stability, the target network is updated less frequently, specifically every $N_{\text{update}}$ steps. This update follows the soft update rule using target update weight $\tau$, where the new target network weights are updated as the weighted average of the policy network weights and the existing target network weights. This gradual update mechanism helps prevent drastic fluctuations and promotes smoother learning. The DQN training procedure is outlined in the pseudocode presented in Algorithm 1.
@@
15:              Set $Q_{\text{target}}=R+\gamma\cdot\max_{a^{\prime}}Q(S_{M}^{\prime},a^{\prime};\theta^{-})$
@@ 5 Computational Study
The warehouse layout and instance generation process employed in this study are derived from Lu et al. (2016). The warehouse configuration consists of a single block with two cross-aisles and 10 parallel aisles, each containing 15 storage locations on either side. Each aisle is 15 meters long, with 3-meter spacing between adjacent aisles. We assume that the depot is located at the end of aisle 6, adjacent to the front cross-aisle, with negligible travel time between these points. The order-picker operates at a speed of 1 meter per second, with a picking time of 5 seconds per item and an additional 1 second for an item drop-off at the depot. The picking device has a capacity ($K$) of 20 items. We assume orders arrive according to a Poisson process and conduct experiments with arrival rates ranging from 0.01 to 0.09 orders per second. For each order arrival rate, we conduct 10 independent simulation runs and report the average results to ensure reliable performance assessment. Each experiment simulates an 8-hour work shift, during which incoming orders are dynamically assigned to the picker to optimize warehouse order-picking throughput. An order is considered complete only after the picker deposits the item at the depot, and not upon retrieval from the respective storage location.

### v2

@@ 4 Methodology
In this section, we present the formulation of the real-time warehouse order picking problem. We model this problem using a finite Markov Decision Process (MDP) with discrete time steps to capture its dynamic nature. Specifically, each time step $t$ involves making a movement or a decision regarding the picker’s actions. At each step, the system state $S_{t}$ is observed, encapsulating information about both the orders, their locations in warehouse and the picker. Based on this state, an action $a_{t}$ is chosen, determining the picker’s movement direction, whether the picker should remain stationary, or if items should be released at the depot. Executing this action leads to a new system state $S_{t+1}$, from which the next action $a_{t+1}$ is determined for the subsequent time step $t+1$. To solve the MDP model, we employ a custom-built Deep Q-learning approach. A high-level overview of this approach is illustrated in Figure 2, with detailed explanations provided in this section.
Figure 2: A visual representation of the proposed solution method
@@ 4.1 MDP Formulation
    State: The system state at time step $t$ is denoted as $\boldsymbol{S_{t}}=(\boldsymbol{S^{p}_{t}},\boldsymbol{S^{o}_{t}})$, where $\boldsymbol{S^{p}_{t}}$ represents the picker state and $\boldsymbol{S^{o}_{t}}$ represents the state of existing orders in the warehouse. In the context of our research, using raw data to represent each state requires managing high-dimensional information. Specifically, representing the picker’s state involves capturing its physical location coordinates on the grid and its current capacity. Similarly, representing existing orders from the raw data requires knowledge of the location coordinates for every storage location and the number of items ordered from each location, which could range from zero to a positive value. Consequently, the dimensionality of the state space derived from raw data can be exceedingly large. To address this complexity and facilitate decision-making in our research, we employ features to characterize the system’s state. These features are designed based on two fundamental assumptions: When the vehicle’s capacity is full, it must immediately return to the depot to drop off the items. If the picker passes a pick location, it will collect all items available there, provided its capacity allows. Additionally, we assume that picking each item requires some time (depending on the application). With this framework in mind, we next present the features used to represent each state.
@@
    The picker state, $\boldsymbol{S^{p}_{t}}=(S^{H}_{t},S^{V_{1}}_{t},S^{V_{2}}_{t},S^{C}_{t})$, consists of four components: $S^{H}_{t}$, $S^{V_{1}}_{t}$, $S^{V_{2}}_{t}$, and $S^{C}_{t}$. Here, $S^{H}_{t}$ represents the horizontal position of the picker, which is the relative position of the picker with respect to cross-aisles, $S^{V_{1}}_{t}$ and $S^{V_{2}}_{t}$ represent the vertical position, which are picker’s positions with respect to pick aisles, and $S^{C}_{t}$ indicates the remaining capacity of the picker. If $S^{C}_{t}=0$, the picker is full and cannot pick up additional items; if $S^{C}_{t}=K$, the picker is empty. The horizontal state $S^{H}_{t}$ can take values of -1, 0, or 1: -1 indicates the picker is at the back cross-aisle, 1 indicates the front cross-aisle, and 0 indicates the picker is within one of the pick aisles. The exact location of the picker within a pick aisle when $S^{H}_{t}=0$ is implicitly encoded in the order state $\boldsymbol{S^{o}_{t}}$. The vertical position is determined by $S^{V_{1}}_{t}$ and $S^{V_{2}}_{t}$, which equal $12n-$ 1 and $2n$ when the picker is at aisle $n$. For example, if the picker is inside aisle $4$ at time point $t$, then the vertical components of the picker state would be $S^{V_{1}}_{t}=7$ and $S^{V_{2}}_{t}=8$. The rationale for representing each pick aisle by two consecutive indices, rather than a single index, will be explained once the order state is described in more details.
@@
    The state of orders, $\boldsymbol{S^{o}_{t}}=(S^{1}_{t},S^{2}_{t},\ldots,S^{2N-1}_{t},S^{2N}_{t})$, where $N$ is the number of pick aisles in the warehouse. Each aisle $n$ is represented by $S^{2n-1}_{t}$ and $S^{2n}_{t}$. These values indicate the state of orders in each pick aisle depending on the picker’s vertical movement direction. $S^{2n-1}_{t}$ represents the potential obtainable reward from aisle $n$ if the picker moves upward in its current aisle, while $S^{2n}_{t}$ represents the potential reward if the picker moves downward. These values are computed by summing the value of each order in the aisle divided by its distance from the picker, based on the direction of movement. The distance between the picker and an order is the smallest number of movements required for picker to reach the order in terms of the number of location storage. These states are calculated as
    $S^{2n-1}_{t}$ $=\displaystyle\sum_{i=1}^{L}\displaystyle\frac{n_{i}}{\rho^{1}_{i}}$ (1)
    $S^{2n}_{t}$ $=\displaystyle\sum_{i=1}^{L}\displaystyle\frac{n_{i}}{\rho^{2}_{i}},$ (2)
@@
    This formulation implicitly captures the importance of the picker’s visiting direction for each aisle. Notably, when $S^{H}_{t}=-1$, then $S^{2n}_{t}=0$ for all aisles since the picker cannot move upward. Similarly, when $S^{H}_{t}=1$, then $S^{2n-1}_{t}=0$ for all aisles. The starting state is $\boldsymbol{S^{o}_{0}}=\boldsymbol{0}$ since there are no orders at time zero, and $\boldsymbol{S^{p}_{t}}=(1,2d-1,2d,K)$, where $d$ is the aisle where depot is located at, and $K$ is the picker capacity. Finally, it is worth mentioning that since the order state $\boldsymbol{S^{o}_{t}}$ is formulated as a vector of size $2N$, we have intentionally chosen the same pattern for vertical components of the picker state for consistency.
@@
    Action: Given the state $S_{t}$, the action $a_{t}$ determines the picker’s decision. The action $a_{t}$ can take on values from 0 to 4. Specifically, $a_{t}=0$ signifies that the picker drops off items if located at the depot, and otherwise results in no movement. Actions with values 1 and 2 correspond to moving to the immediate next aisle, specifically to the right and left aisles, respectively. In this study, actions are constrained to feasible options only. For instance, when the picker is situated within a pick aisle (i.e., $S^{H}_{t}=0$), actions 1 and 2 are disallowed. Actions 3 and 4 correspond to moving the picker upward and downward, respectively. The movement associated with Actions 3 and 4 is not necessarily limited to one unit; it should be sufficient to alter the state of the system as defined in the state transition explanation. Additionally, when the picker is in the back cross-aisle (i.e., $S^{H}_{t}=-1$), action 3 is prohibited, and when the picker is in the front cross-aisle (i.e., $S^{H}_{t}=1$), action 4 is prohibited.
@@
    $r_{t}=\begin{cases}-1&\text{if }a_{t}=0\text{ and picker not at the depot},\\ R\times(K-S^{C}_{t})\times\alpha&\text{if }a_{t}=0\text{ and picker at the depot},\\ -n^{m}_{t}+R\times n^{p}_{t}&\text{otherwise}.\end{cases}$ (3)
    The components of the reward function are interpreted as follows:
    1. (a)
      Case: $a_{t}=0$ and the picker is not at the depot. This action signifies that the picker remains stationary at its current location. A negative reward of $-1$ is assigned in this case, incentivizing the picker to move when a potential positive reward can be obtained rather than staying idle. This penalty aligns with the penalty for moving one storage location, thereby encouraging exploration without overly penalizing idle actions. By making the penalty for staying still equal to the cost of moving, the policy is encouraged to weigh the potential rewards of movement more effectively.
    2. (b)
      Case: $a_{t}=0$ and the picker is at the depot. This action represents unloading all items from the picker. The reward for this action is proportional to the number of items being unloaded, which is given by $K-S^{C}_{t}$, where $S^{C}_{t}$ denotes the current capacity of the picker. The parameter $\alpha$ modulates the reward for unloading items relative to picking them up, with its value ranging between 0 and 1. Smaller values of $\alpha$ prioritize picking up items over unloading, encouraging the picker to maximize its capacity before heading to the depot, which helps reduce travel distance. Conversely, larger values of $\alpha$ balance the value of unloading and picking, which minimizes both waiting and delivery times. For example, suppose $R=8$, $K=10$, $S^{C}_{t}=7$, and $\alpha=0.5$. In this case, the picker unloads $3$ items at the depot, receiving a reward of:
      $R\times(K-S^{C}_{t})\times\alpha=8\times 3\times 0.5=12.$
    3. (c)
      Case: $a_{t}\neq 0$. Any action other than 00 leads to the picker moving to a new location. The reward in this case depends on both the distance traveled (number of storage locations) and the number of items picked up during this movement.
      • The number of storage locations traversed is penalized by a term $n^{m}_{t}$, representing the effort or cost of movement.
      • If items are picked up along the way, the reward is augmented by a term $R\times n^{p}_{t}$, where $n^{p}_{t}$ is the number of items picked up.
      • For actions $a_{t}=1$ and $a_{t}=2$, which correspond to moving left and right within the cross aisles, no items are picked up. Hence, the reward includes only the penalty for movement ($-n^{m}_{t}$).
      • For actions $a_{t}=3$ and $a_{t}=4$, which involve moving and picking up orders, the reward reflects both penalties for movement and the reward for picking items. For example, suppose the picker moves 7 storage locations and picks up 2 items, with $R=8$. The reward is calculated as:
        $-n^{m}_{t}+R\times n^{p}_{t}=-7+8\times 2=9.$
Additional discussion regarding the rationale for the feature-based state representation proposed in this study is provided in A.
Figure 3: Proposed Deep Neural Network Architecture
@@ 4.2 Proposed Approach
Our objective is to train a policy that aims to maximize the discounted cumulative reward $R_{t}=\sum_{k=0}^{\infty}\gamma^{k}\times r_{t+k}$, where $R_{t}$ is referred to as the return, and $10<\gamma<$ 1 is the discount factor that balances the importance between immediate and future rewards. Small values of $\gamma$ indicate that immediate actions are more significant than future actions, while larger values of $\gamma$ place greater emphasis on future rewards.
The fundamental concept behind Q-learning (Watkins, 1989) is that if we had access to a function $Q^{*}:\text{State}\times\text{Action}\rightarrow\mathbb{R}$, which could inform us of the expected return if we were to take a particular action in a given state, we could straightforwardly construct a policy that maximizes our rewards: $\pi^{*}=\arg\max_{a}Q^{*}(s,a)$, where $Q^{*}(s,a)=\max(E_{\pi^{*}}[R_{t}\mid s_{t}=s,a_{t}=a])$. However, due to the complexity of the environment, we do not have direct access to $Q^{*}$. Nevertheless, since neural networks are universal function approximators, we can construct a neural network and train it to approximate $Q^{*}$.
@@ 4.2.1 Architecture of Deep Neural Network
The deep neural network architecture designed for this study is specifically tailored to handle the state representation and predict optimal actions for the picker in a dynamic warehouse environment. The network serves as a Q-network, approximating the action-value function $Q(s,a)$, and consists of several fully connected layers to process both the picker state and the order state separately before combining them for the final action prediction. Figure 3 provides a summary of the architecture and it is structured as follows:
  • Input Layers: The input to the neural network is divided into two parts: the picker state $S^{p}_{t}$ and the order state $S^{o}_{t}$. The picker state $S^{p}_{t}$ is a four-dimensional vector represented as $(S^{H}_{t},S^{V_{1}}_{t},S^{V_{2}}_{t},S^{C}_{t})$, while the order state $S^{o}_{t}$ is a vector of size $2N$, where $N$ is the number of aisles in the warehouse. Note that the input to the neural network is the state representation as defined earlier, comprising both picker state ($S^{p}_{t}$) and order state ($S^{o}_{t}$). This ensures that the neural network processes the exact state information defined for the problem.
@@
We utilize this deep neural network architecture to capture the complex interactions between the picker and the orders in the warehouse, enabling it to predict optimal actions that maximize the cumulative reward over time.
4.2.2 Comparison with Alternative Neural Network Architectures
To illustrate why a fully connected feed-forward neural network (FCN) is particularly well-suited to our problem setting, we compare it with other common neural network architectures. Table 2 summarizes key architectural choices, their typical computational complexity, and their applicability to the warehouse order-picking domain.
Table 2: Comparison of common neural network architectures
Architecture
Computational Complexity
Suitability
Fully Connected
Network (FCN)
(Our Approach)
$\mathcal{O}(n\cdot m)$ per layer,
where $n$ = # input features,
$m$ = # outputs.
Well-suited for structured,
tabular-like input
Convolutional Neural
Network (CNN)
$\mathcal{O}(H\cdot W)$ per layer,
where $H\times W$ = 2D input size.
Assumes fixed kernel and
channel sizes.
Ideal for spatial/
image-based data
Graph Neural
Network (GNN)
$\mathcal{O}(|V|+|E|)$ per layer,
where $|V|$ = # nodes, $|E|$ = # edges.
Assumes sparse graph and fixed
feature dimensions.
Useful for explicit graph
representations
Recurrent Neural
Network (RNN)
$\mathcal{O}(T)$ per layer,
where $T$ = sequence length.
Assumes fixed hidden and input
dimensions across steps.
Effective for temporal
or sequential data
Given the aggregated vector-based state representation introduced in Section 4.1, the proposed framework utilizes two types of inputs: a relatively low-dimensional picker state vector $(S^{H}_{t},S^{V_{1}}_{t},S^{V_{2}}_{t},S^{C}_{t})$ and a higher-dimensional, yet still tabular, order state vector $(S^{1}_{t},S^{2}_{t},\ldots,S^{2N}_{t})$, which captures aggregated aisle-level demand. These inputs constitute structured tabular data, as opposed to raw image-based inputs or explicit spatial grids. Accordingly, a fully connected network (FCN) is adopted, given its effectiveness in processing tabular data with relatively low computational overhead. This makes the architecture well-suited for dynamic and time-sensitive tasks such as real-time order-picking decisions. Moreover, spatial relationships and capacity constraints are already accounted for through targeted feature engineering—such as aisle-level demand aggregation and distance-based reward shaping—thereby mitigating the need for more specialized architectures like convolutional neural networks (CNNs), graph neural networks (GNNs), or recurrent neural networks (RNNs). These alternatives are typically employed when handling data with strong spatial, graphical, or temporal dependencies, which are not explicitly present in our setting.
@@ 4.2.3 Training of Deep Neural Network
15:              Set $Q_{\text{target}}=R+\gamma\times\max_{a^{\prime}}Q(S_{M}^{\prime},a^{\prime};\theta^{-})$
@@
The training procedure for each mini-batch is a multi-step process. First, the states, next states, actions, and rewards within the batch are stacked for parallelization. Then, the policy network is set to training mode, while the target network enters evaluation mode. This setup allows for the computation of state-action values using the policy network, and expected state-action values using the target network. The discrepancy between these values is then used to calculate the loss, and we utilize the Huber loss function (Huber, 1992) for robustness. The calculated loss is then backpropagated through the policy network, facilitating its optimization. To ensure stability, the target network is updated less frequently, specifically every $N_{\text{update}}$ steps. This update follows the soft update rule using target update weight $\tau$, where the new target network weights are updated as the weighted average of the policy network weights and the existing target network weights. This gradual update mechanism helps prevent drastic fluctuations and promotes smoother learning. The Deep Q-Network (DQN) training procedure is outlined in the pseudocode presented in Algorithm 1.
@@ 5 Computational Study
The warehouse layout and instance generation process employed in this study are derived from Lu et al. (2016). The warehouse configuration consists of a single block with two cross-aisles and 10 parallel aisles, each containing 15 storage locations on either side. Each aisle is 15 meters long, with 3-meter spacing between adjacent aisles. We assume that the depot is located at the end of aisle 6, adjacent to the front cross-aisle, with negligible travel time between these points. The order-picker operates at a speed of 1 meter per second, with a picking time of 5 seconds per item and an additional 1 second for an item drop-off at the depot. The picking device has a capacity ($K$) of 20 items. We assume that orders arrive following a Poisson process and are distributed uniformly throughout the warehouse. We conduct experiments with arrival rates ranging from 0.01 to 0.09 orders per second. For each order arrival rate, we conduct 10 independent simulation runs and report the average results to ensure reliable performance assessment. Each experiment simulates an 8-hour work shift, during which incoming orders are dynamically assigned to the picker to optimize warehouse order-picking throughput. An order is considered complete only after the picker deposits the item at the depot, and not upon retrieval from the respective storage location.